A store has candy worth $0.90 a pound and candy worth $1.20 a pound. if the owners want 60 pounds of candy worth $1 a pound, how many pounds of each candy should they use?
Answer by richwmiller(17219)

.9x+1.2y=60
x+y=60
.9x+1.2y=x+y
.1x=.2y
x=2y
2y+y=60
3y=60
y=20 at $.90
x=40 at $1.20
check
.9*40+1.2*20=60
36+24=60
60=60
ok